OpenStudy (anonymous):

Let u = <7, -3>, v = <-9, 5>. Find 4u - 3v.

OpenStudy (legends):

I don't get it @Data_LG2

OpenStudy (anonymous):

Hi Welcome to Openstudy! \(\sf 4u - 3v = 4 (7, -3) - 3 (-9, 5)\\ \hspace{25 pt} =4(7)i - 3(-9)i + 4(-3)j-3(5)j\) simplify